a certain game consist of rolling a single fair die and pays off as follows: $5 for a $6, $4 for a 5, $3 for a 4 and no payoff otherwise. find the expected winnings for this game
Answer #1

SOLUTION :

  • Let X is a random variable shows the winning amount. Here X can take values $5, $4, $3 and $0.

    When a single fair die is rolled the possible outcomes are 1, 2, 3, 4, 5, and 6.

    Out of 6 outcomes, 1 is 6 so when X= $5 then

    P(X=$5) = 1/6

    Out of 6 outcomes, 1 is 5 so when X= $4 then

    P(X=$4) = 1/6

    Out of 6 outcomes, 1 is 4 so when X= $3 then

    P(X=$3) = 1/6

    And

    P(X=$0) = 1- P(X=$5) - P(X=$4) - P(X=$3) = 3/6

    Following table shows calculations for expected winning:

    X P(X=x) xP(X=x)
    5 1/6 5/6
    4 1/6 4/6
    3 1/6 3/6
    0 3/6 0
    Total 2

So expected winning is

​​​​​​​

NOTE : IF YOU SEE MY ANSWER PLEASE RATE MY ANSWER FIRST.THANKS.